### v1

@@ Abstract
3D visual grounding aims to identify the target object within a 3D point cloud scene referred to by a natural language description. While previous works attempt to exploit the verbo-visual relation with proposed cross-modal transformers, unstructured natural utterances and scattered objects might lead to undesirable performances. In this paper, we introduce DOrA, a novel 3D visual grounding framework with Order-Aware referring. DOrA is designed to leverage Large Language Models (LLMs) to parse language description, suggesting a referential order of anchor objects. Such ordered anchor objects allow DOrA to update visual features and locate the target object during the grounding process. Experimental results on the NR3D and ScanRefer datasets demonstrate our superiority in both low-resource and full-data scenarios. In particular, DOrA surpasses current state-of-the-art frameworks by 9.3% and 7.8% grounding accuracy under 1% data and 10% data settings, respectively.
Keywords:
3D Computer Vision 3D Visual Grounding
@@ 1 Introduction
As pioneers of 3D visual grounding, Referit3D [2] and ScanRefer [8] are two benchmark datasets that build upon the point cloud scene provided by the ScanNet [12] dataset. The former presents a graph neural network (GNN) [36]-based framework to explicitly learn the object spatial relations as the baseline. The latter designs a verbo-visual cross-modal feature extraction and fusion pipeline as the baseline. Several subsequent methods are presented [20, 46, 4, 31, 21, 15, 19, 11, 40], which either augment the visual features [20, 46, 4, 31], utilize specialized cross-modal feature alignment frameworks [21], or exploit spatial relations with designed frameworks/modules [17, 15, 19, 11, 40] are proposed to solve the 3D visual grounding problem. However, these approaches use a referring head to localize the target object directly. Without explicitly considering any additional information about the anchor objects mentioned in the description, models must implicitly discover the relation between the anchor objects and the target object. They may be misled by other similar objects presented, as pointed out in [5]. To overcome this issue, some approaches [49, 1, 42] propose to incorporate anchor objects during training by including their label annotations [18, 5]. Nonetheless, human annotators are usually required to obtain this additional linguistic information [49, 1], causing potential difficulty in scaling up to larger datasets for real-world applications.
@@
To unload the need for human annotators, recent approaches [18, 5] leverage large language models (LLMs) for automatic dissection of the descriptions and generation of prior linguistic knowledge. For example, NS3D [18] utilizes Codex [10] to parse descriptions into nested expressions and designs a neuro-symbolic framework to find the target object step-by-step. However, NS3D is restricted to fixed-template relations between objects (e.g., below/above, near/far, etc.) and cannot be easily extended to arbitrary natural descriptions. Inspired by the mechanism of human perception system [32, 9], CoT3DRef [5] generates the referential order of a description that points from anchor objects to the final target object. For example, for a description “Find the water bottle on the table nearest to the door.”, the referential order is generated as { “door” (anchor), “table” (anchor), “water bottle” (target) }. A rule-based algorithm is presented to establish matching between each class name in the referential order and a corresponding anchor/target object in the scene at once, with the matching results being used to guide a transformer-based module to predict the final target object. However, if the matching results calculated by the rule-based method are not accurate, the performance might be degraded.
@@
In this paper, we propose a 3D Visual Grounding Framework with Order-Aware Referring (DOrA). By utilizing a Large Language Model (LLM) of GPT-3.5-Turbo [48], DOrA exploits the awareness of anchor objects with the associated referential order from the textual description, as depicted in LABEL:fig:teaser. With this referential order of anchor objects as guidance, a series of Object Referring blocks is deployed in DOrA to process the extracted anchor/target objects, each performing feature enhancement to update the visual features of corresponding objects. Since only the ground-truth identity of the target object is available during training (no ground-truth referential order observed), we additionally introduce a pre-training strategy to augment accurate labels and referential orders for anchor/target objects. By conducting extensive experiments on real-world benchmark datasets, we show that our DOrA is comparable to current state-of-the-art 3D visual-grounding methods on the standard setting and further outperforms them significantly when training data is limited.
@@
  • We present a 3D Visual Grounding Framework with Order-Aware Referring (DOrA) for 3D visual grounding via natural descriptions.
  • With a two-stage in-context learning technique on LLM, DOrA infers a series of anchor/target objects in a referential order to guide the grounding process.
  • DOrA deploys a series of Object-Referring blocks that follows the referential order and progressively refines the features of anchor/target objects for improved grounding performances.
  • A pre-training scheme is introduced for DOrA, which augments reliable labels and accurate referential orders of anchor/target objects as additional training examples, ensuring the awareness of the identities and referential orders of anchor objects.
@@ 2 Related Work
2D visual grounding aims to locate the target object in an image referred to by a natural language description, with various approaches being proposed in recent years [23, 28, 51, 29, 47, 39, 38, 45]. Among them, verbo-visual feature alignment frameworks [23, 28, 51, 29] have proven themselves to be an effective way to equip models with abilities of tackling description contexts and image semantics simultaneously. Particularly, as one of the pioneers, MDETR [23] extends DETR [7], an end-to-end object detection framework, to incorporate text modalities with the proposed text-image alignment contrastive losses. GLIP [51] takes a step forward to improve the performance of visual grounding by proposing unified multi-task learning that includes object localization and scene understanding tasks, showing that these tasks could gain mutual benefits from each other. Grounding DINO [29] further designs the large-scale grounding pretraining for DINO [16], reaching the capability of open-set grounding. Although great progress is achieved, extending these 2D visual grounding methods to 3D scenarios is not easy due to the additional depth information in 3D data that triggers more complicated object arrangements and more complex descriptions to describe the relations between objects, leaving 3D visual grounding as an unsolved research area.
@@
To better identify the target object of interest, some [46, 4, 31, 20] intend to produce richer input semantic information for the model. For example, [46, 4, 31] introduce image features by acquiring 2D images of the scene to obtain more color/shape information. On the other hand, MVT [20] projects the point cloud into multiple views for more position information. Although these approaches have achieved great progress in dealing with scattered object arrangements, such methods typically extract a global, sentence-level feature [13, 30] from the given natural description. As a result, detailed information within the description (e.g., the target object, anchor objects, and their relations) may not be preserved and leveraged properly, potentially reducing training efficiency and prediction accuracy of the model as discussed in [5].
@@
To address the above issue, some works have put their efforts into mining the natural descriptions to acquire additional prior knowledge for more efficient and effective learning [49, 1, 42, 18, 5]. Specifically, ScanEnts3d [1] and 3DPAG [49] recruit human annotators to establish one-to-one matching between each anchor object mentioned in the description and the corresponding object entity in the 3D scene. With this additional information, they design dense word-object alignment losses to improve the training. However, annotating one-to-one correlation requires considerable labor. For example, it takes 3664 hours of workforce commitment in 3DPAG to annotate anchor objects for 88k descriptions. To eliminate the need for human annotators, NS3D [18] makes the first attempt to use the LLM. It leverages Codex [10] to parse fixed-template descriptions into nested logical expressions, followed by a neural-symbolic framework to execute the logical expressions implemented as programmatic functions.
@@
Unfortunately, NS3D cannot handle arbitrary natural descriptions well, which might generate complicated expressions and unforeseen functions, hindering the framework from successful execution [18]. Recently, CoT3DRef [5] proposes to acquire the referential order of the description, listing from anchor objects to the final target object, using LLMs. Also, the authors design a rule-based searching method using a traditional sentence parser [37] to construct the one-to-one matching between class names in the order and potential anchor/target objects in the scene at once. The matching information is encoded by positional encoding as additional inputs for the proposed transformer-based CoT module to learn the referring process implicitly and predict the final target object. Nevertheless, since rule-based matching is not guaranteed to be correct, such additional inputs might result in noisy information and degrade the grounding process, impeding correct target prediction and affecting training stability.
@@ 3.1.2 Framework overview
As shown in Fig. 1, DOrA is composed of $B$ consecutive Object-Referring blocks $\{R_{1},\cdots,R_{B}\}$ to progressively locate the target object. In particular, by taking both $P$ and $D$ as the inputs, DOrA utilizes the Object-Referring blocks $R_{1:B}$ in Fig. 1 to sequentially produce anchor objects to guide the grounding process. Each $R_{i}$ takes the object feature $F_{i}\in\mathbb{R}^{K\times d_{i}}$ and the text feature $T\in\mathbb{R}^{(|D|+1)\times 768}$ as the inputs. Note that $T$ contains a ${1\times 768}$-dimensional sentence-level feature and ${|D|\times 768}$-dimensional word-level feature, where $|D|$ denotes the length of $D$ after tokenization. By observing $F_{i}$ and $T$, the Object-Referring block $R_{i}$ aims to produce the refined feature $F_{i+1}$ along with the updated anchor/target objects and their relations for grounding purposes.
@@
To enable our Object-Referring blocks to capture proper information about the anchor/target objects, we apply a Large Language Model (LLM) to $D$ to generate a Referential Order $O_{1:B}=\{O_{1},\cdots,O_{B}\}$ that mimics human perception system of searching target object [32, 9] by extracting and arranging the class names of the anchor and target objects. Specifically, $\{O_{1},\cdots,O_{B-1}\}$ represent the class names of the anchor objects, and $O_{B}$ is the class name of the target object. Note that $O_{i:B}$ is observed by the $i$-th Object-Referring block $R_{i}$ as guidance, which updates the features of anchor/target objects with the proposed Feature Enhancing (FE) module. Since there is no ground truth referential order of the produced anchor objects, we also introduce a pre-training strategy for training DOrA. This is achieved by synthesizing accurate referential order and anchor/target object labels. With the above design, DOrA can be applied for 3D grounding tasks and achieve satisfactory performances. We now detail the design of our DOrA in the following subsections.
Figure 1: Architecture of our 3D Visual Grounding Framework with Order-Aware Referring (DOrA). By taking a point cloud scene $C$ and a natural description $D$ as inputs, our DOrA produces a referential order of anchor objects $O_{1:B}$ and conduct Object-Referring blocks $R_{1:B}$ to locate the target object progressively.
3.2 Referential Ordering using LLM
To have Object-Referring blocks $\{R_{1},\cdots,R_{B}\}$ to locate the target object properly, it is desirable to extract a proper referential order $O_{1:B}$ from the input description $D$. With such a referring path constructed, visual features of the associated objects can be updated for grounding purposes. This is inspired by the idea of Chain-of-Thoughts [41, 26] in LLM, as noted in CoT3DRef [5]. To achieve such a parsing task, we apply GPT-3.5-Turbo [48] as the description parser using in-context learning (ICL) [14], as depicted in Fig. 1(a).
However, it is not trivial to have LLM output a referential order from $D$ due to lengthy and noisy descriptions. For example, for an input description “Look at the king-size bed in the room next to a green chair, Find the pillow on the bed. Not the pillow on the sofa which is also beside the chair.”, one would expect first to find the green chair then the king-size bed next to the chair, and finally, the pillow on the bed. Therefore, the ideal referential order is {“chair”, “bed”, “pillow”}. In the above example, the sentence “Not the pillow on the sofa which is also beside the chair.” is redundant since one can find the correct target object without this information. If we apply the LLM directly to the original description, the model may be misled by this redundant information and generate a referential order containing the object “sofa”.
To tackle the above problem, we conduct a two-stage in-context learning (ICL) scheme to remove redundant information in $D$ before producing $O_{1:B}$, as depicted in Fig. 1(a). With the given target object $O_{B}$, we predict a summarized description $D^{{}^{\prime}}$ to remove redundant information in $D$ in the first stage. Then, the entire referential order $O_{1:B}$ given $D^{{}^{\prime}}$ and $O_{B}$ is produced in the second stage. For each stage of our ICL, 10 examples are provided as demonstrations of the input prompts for LLM to predict $O_{1:B}$. Due to page limitations, demonstration examples and the complete prefix prompts are presented in the supplementary materials. With $O_{1:B}$ obtained, the order-aware object referring process can be processed accordingly, as we detail next.
(a) Referential order generation via LLM. A two-stage ICL is deployed to remove redundant information from $D$, forming a descriptive order $O_{1:B}$ for locating the target object $O_{B}$.
@@ 3.2 Referential Ordering using LLM
(b) Feature Enhancement (FE). Taking $R_{1}$ as an example, FE processes the objects described mentioned in $O_{1:B}$ and the relations between them by attending the masked feature $F^{mask}_{1}$. Thus, only object features related to $O_{1:B}$ would be refined as $F_{2}$.
Figure 2: Illustration of referential order generation and feature enhancement in DOrA.
@@
Given the object proposals $P$, the corresponding labels $L=\{l_{1},\cdots,l_{K}\}$, the encoded text features $T$, and the derived referential order $O_{1:B}$ the as inputs, our DOrA deploys a series of Object-Referring blocks $\{R_{1},\cdots,R_{B}\}$ to perform the grounding task. As depicted in Fig. 1, this referring process is conducted by leaving out an anchor object and updating the visual features in each step, until only the final step locates the target object of interest. Thus, the deployment of Object-Referring blocks allows one to focus on the anchor/target objects, so that their visual features and spatial relations between them can be exploited with those of irrelevant objects disregarded.
Take the $i$-th referring block observing $O_{i:B}$ for example, a masked feature $F^{mask}_{i}=F_{i}\odot M_{i}$ is derived by applying a Hadamard product between $F_{i}$ and a $K$-dimensional binary mask $M_{i}$ to replace features of object proposals in $F_{i}$ not belonging to any of the object classes in $O_{i:B}$. Such a masking strategy ensures that $F^{mask}_{i}$ contains objects described in $O_{i:B}$ and hence explicitly suppresses effects of irrelevant objects that are not in our interests. Thus, the $j$-th entry of $M_{i}$ (denoted as $m_{ij}$) is defined as:
@@
3.3.2 Object feature enhancement.

### v2

@@ Abstract
3D visual grounding aims to identify the target object within a 3D point cloud scene referred to by a natural language description. Previous works usually require significant data relating to point color and their descriptions to exploit the corresponding complicated verbo-visual relations. In our work, we introduce Vigor, a novel Data-Efficient 3D Visual Grounding framework via Order-aware Referring. Vigor leverages LLM to produce a desirable referential order from the input description for 3D visual grounding. With the proposed stacked object-referring blocks, the predicted anchor objects in the above order allow one to locate the target object progressively without supervision on the identities of anchor objects or exact relations between anchor/target objects. In addition, we present an order-aware warm-up training strategy, which augments referential orders for pre-training the visual grounding framework. This allows us to better capture the complex verbo-visual relations and benefit the desirable data-efficient learning scheme. Experimental results on the NR3D and ScanRefer datasets demonstrate our superiority in low-resource scenarios. In particular, Vigor surpasses current state-of-the-art frameworks by 9.3% and 7.6% grounding accuracy under 1% data and 10% data settings on the NR3D dataset, respectively.
@@ 1 Introduction
As pioneers of 3D visual grounding, Referit3D [2] and ScanRefer [8] are two benchmark datasets that build upon the point cloud scene provided by the ScanNet [12] dataset. The former presents a graph neural network (GNN) [36]-based framework to explicitly learn the object spatial relations as the baseline. The latter designs a verbo-visual cross-modal feature extraction and fusion pipeline as the baseline. Following the settings of the aforementioned benchmarks, several subsequent methods are presented [20, 46, 4, 31, 21, 15, 19, 11, 40]. However, these approaches use a referring head to localize the target object directly. Without explicitly considering any additional information about the anchor objects mentioned in the description, models must implicitly discover the relation between the anchor objects and the target object. They may be misled by other similar objects presented, as pointed out in [5]. To overcome this issue, some approaches [50, 1, 42] propose to incorporate anchor objects during training by including their label annotations [18, 5]. Nonetheless, human annotators are usually required to obtain this additional linguistic information [50, 1], causing potential difficulty in scaling up to larger datasets for real-world applications.
Figure 1: Referential orders for 3D grounding. The order manifests an anchor-to-target referring process that helps the grounding model identify the target object described in the input.
@@
To eliminate the need of human annotation for training grounding models, recent approaches [18, 5] leverage large language models (LLMs) for automatic dissection of the descriptions and generation of prior linguistic knowledge. For example, NS3D [18] utilizes Codex [10] to parse descriptions into nested expressions and designs a neuro-symbolic framework to find the target object step-by-step. However, it only considers fixed-template relations between objects (e.g., below/above, near/far, etc.) and cannot be easily extended to arbitrary natural descriptions. Inspired by the mechanism of human perception system [32, 9], CoT3DRef [5] generates the referential order of a description that points from anchor objects to the final target object using LLM. For example, for a description “Find the water bottle on the table nearest to the door.”, the referential order is generated as { “door” (anchor), “table” (anchor), “water bottle” (target) }. Additionally, it utilizes a rule-based algorithm to localize the identities of the above anchor/target objects, which guides a transformer-based module to predict the final target object. However, as noted in [5], such rule-based identity prediction might not be applicable for scenarios with complex language descriptions.
@@
In this paper, we propose a data-efficient 3D Visual Grounding framework via Order-aware Referring (Vigor). Leveraging the LLM-parsed referential order, Vigor exploits the awareness of anchor objects from the textual description, as depicted in Fig. 1. With such ordered anchor objects as guidance, a series of Object Referring blocks are deployed to process the corresponding objects, each performing feature enhancement to update the visual features of corresponding objects. Since only the ground-truth grounding information of the target object is available during training (no ground-truth referential order observed), we additionally introduce a unique warm-up learning strategy to Vigor. This pre-training scheme can be viewed as augmenting object labels and referential orders to initialize Vigor so that it can be realized in data-efficient training schemes. Our experiments on real-world benchmark datasets confirm that Vigor performs favorably against recent 3D visual grounding methods, especially when the size of training data is limited.
@@
  • We present a Data-Efficient 3D Visual Grounding Framework via Order-Aware Referring (Vigor), which performs 3D visual grounding from natural description inputs.
  • By utilizing sequential yet consecutive Object Referring blocks, Vigor is able to locate anchor/target objects mentioned in the description by considering plausible referential orders established by LLM.
  • We introduce a warm-up strategy that introduces the model with the ability to locate anchor/target object identities by synthesizing training examples of reliable labels and referential orders.
  • Through comprehensive experiments, we show that Vigor achieves satisfactory performances in various low-source settings, surpassing current grounding approaches significantly.
@@ 2 Related Work
2D visual grounding aims to locate the target object in an image referred to by a natural language description, with various approaches being proposed in recent years [23, 28, 51, 29, 47, 39, 38, 45]. Among them, verbo-visual feature alignment frameworks [23, 28, 51, 29] have proven themselves to be an effective way to equip models with abilities to tackle description contexts and image semantics simultaneously. Particularly, as one of the pioneers, MDETR [23] extends DETR [7], an end-to-end object detection framework, to incorporate text modalities with the proposed text-image alignment contrastive losses. GLIP [51] takes a step forward to improve the performance of visual grounding by proposing unified multi-task learning that includes object localization and scene understanding tasks, showing that these tasks could gain mutual benefits from each other. Grounding DINO [29] further designs the large-scale grounding pretraining for DINO [16], reaching the capability of open-set grounding. Although great progress is achieved, extending these 2D visual grounding methods to 3D scenarios is not easy due to the additional depth information in 3D data that triggers more complicated object arrangements and more complex descriptions to describe the relations between objects, leaving 3D visual grounding as an unsolved research area.
@@
To better identify the target object, [46, 4, 31, 20] aim to produce richer input semantic information for learning the grounding model. For example, [46, 4, 31] introduce image features by acquiring 2D images of the scene to obtain more color/shape information. MVT [20] projects the point cloud into multiple views for more position information. Although these approaches have achieved great progress in dealing with scattered object arrangements, such methods typically extract a global, sentence-level feature [13, 30] from the given natural description. As a result, detailed information such as the target object, anchor objects, and their relations may not be preserved and leveraged properly, potentially reducing training efficiency and prediction accuracy as discussed in [5].
@@
To address the above issue, some works have put their efforts into mining the natural descriptions to acquire additional prior knowledge for improved learning [50, 1, 42, 18, 5]. Specifically, ScanEnts3d [1] and 3DPAG [50] recruit human annotators to establish one-to-one matching between each anchor object mentioned in the description and the corresponding object entity in the 3D scene. With such additional information, they design dense word-object alignment losses to improve the training. However, annotating one-to-one text-3D relations requires considerable labor effort. For example, it takes more than 3600 hours of workforce commitment in 3DPAG to annotate anchor objects for 88k descriptions.
2.3 Data-Efficient 3D Visual Grounding
To eliminate the need for human annotators for learning grounding models, NS3D [18] makes the first attempt to use the LLM. It leverages Codex [10] to parse fixed-template descriptions into nested logical expressions, followed by a neural-symbolic framework to execute the logical expressions implemented as programmatic functions. By doing so, NS3D correctly locates the anchor/target objects mentioned in the given description and achieves impressive performance with only $0.5\%$ training data on synthetic datasets.
@@ 2.3 Data-Efficient 3D Visual Grounding
Unfortunately, NS3D is not designed to handle arbitrary natural descriptions, resulting in complicated expressions and unforeseen functions and hindering the framework from successful execution [18]. Recently, CoT3DRef [5] proposes to utilize LLMs to acquire the referential order of the description, listing from anchor objects to the final target object. It deploys a rule-based searching method using a traditional sentence parser [37] to construct the one-to-one matching between class names in the order and potential anchor/target objects in the scene at once. The matching information is encoded by positional encoding as pseudo labels of anchor/target objects and as additional inputs for the proposed transformer-based CoT module to learn the referring process implicitly and predict the final target object. By the above design, CoT3DRef is able to reduce the training data to $10\%$ while preserving competitive performance on real-world datasets. Nevertheless, since the parsed pseudo labels may not be accurate, taking them as additional inputs might result in noisy information and degrade the grounding process, impeding correct target prediction and affecting training stability.
@@ 3 Methodology
Figure 2: Architecture of our 3D Visual Grounding Framework with Order-Aware Referring (Vigor). By taking a point cloud scene $C$ and a natural description $D$ as inputs, our Vigor produces a referential order of anchor/target objects $O_{1:B}$ and conduct Object-Referring blocks $R_{1:B}$ to locate the target object progressively.
@@ 3.1.2 Framework overview
As shown in Fig. 2, Vigor is composed of $B$ consecutive Object-Referring blocks $\{R_{1},\cdots,R_{B}\}$ to progressively locate the target object. In particular, by taking both $P$ and $D$ as the inputs, Vigor utilizes the Object-Referring blocks $R_{1:B}$ in Fig. 2 to sequentially produce anchor objects to guide the grounding process. Each $R_{i}$ takes the object feature $F_{i}\in\mathbb{R}^{K\times d_{i}}$ and the text feature $T\in\mathbb{R}^{(|D|+1)\times 768}$ as the inputs. Note that $T$ contains a ${1\times 768}$-dimensional sentence-level feature and ${|D|\times 768}$-dimensional word-level feature, where $|D|$ denotes the length of $D$ after tokenization. By observing $F_{i}$ and $T$, the Object-Referring block $R_{i}$ aims to produce the refined feature $F_{i+1}$ along with the updated anchor/target objects and their relations for grounding purposes.
@@
To enable our Object-Referring blocks to capture proper information about the anchor/target objects, we apply a Large Language Model (LLM) to $D$ to generate a Referential Order $O_{1:B}=\{O_{1},\cdots,O_{B}\}$ that mimics human perception system of searching target object [32, 9] by extracting and arranging the class names of the anchor and target objects, similar to [5]. Specifically, $\{O_{1},\cdots,O_{B-1}\}$ represent the class names of the anchor objects, and $O_{B}$ is the class name of the target object (please refer to Supp. F for details of Referential Order generation). Note that $O_{i:B}$ is observed by the $i$-th Object-Referring block $R_{i}$ as guidance, which updates the features of anchor/target objects with the proposed Feature Enhancing (FE) module. Since the ground truth referential order is not available during training, we introduce a unique warm-up strategy for training Vigor. This is achieved by synthesizing accurate referential order and anchor/target object labels. It is worth noting that, with the above design, Vigor can be applied for 3D grounding tasks and achieve satisfactory performances with a respectively limited amount of training data. We now detail the design of our Vigor in the following subsections.
@@
(a) Feature Enhancement (FE). Taking $R_{1}$ as an example, FE processes the objects described mentioned in $O_{1:B}$ and the relations between them by attending the masked feature $F^{mask}_{1}$. Thus, only object features related to $O_{1:B}$ would be refined as $F_{2}$.
(b) Synthesizing a referential order and description for order-aware learning warm-up. Given $P$ and $L$, several class names are sampled to construct $D^{aug}$. The identities of anchor/target objects $p^{aug}_{1:B}$ described in $D^{aug}$ are then located by considering the center coordinates and class name of each proposal. By the above design, the augmented referential order $O^{aug}_{1:B}$ is uniquely determined (i.e., the appearance order of each sampled class name in $D^{aug}$).
Figure 3: Illustration of feature enhancement and synthesizing warmup data in Vigor.
@@
Given the object proposals $P$, the corresponding labels $L=\{l_{1},\cdots,l_{K}\}$, the encoded text features $T$, and the derived referential order $O_{1:B}$ the as inputs, our Vigor deploys a series of Object-Referring blocks $\{R_{1},\cdots,R_{B}\}$ to perform the grounding task. As depicted in Fig. 2, this referring process is conducted by leaving out an anchor object and updating the visual features in each step until only the final step locates the target object of interest. Thus, the deployment of Object-Referring blocks allows one to focus on the anchor/target objects so that their visual features and spatial relations between them can be exploited while those of irrelevant objects are disregarded.
Take the $i$-th referring block observing $O_{i:B}$ for example, a masked feature $F^{mask}_{i}=F_{i}\odot M_{i}$ is derived by applying a Hadamard product between $F_{i}$ and a $K$-dimensional binary mask $M_{i}$ to replace features of object proposals in $F_{i}$ not belonging to any of the object classes in $O_{i:B}$. Such a masking strategy ensures that $F^{mask}_{i}$ contains objects described in $O_{i:B}$ and hence explicitly suppresses the effects of irrelevant objects that are not in our interests. Thus, the $j$-th entry of $M_{i}$ (denoted as $m_{ij}$) is defined as:
@@
3.2.2 Object feature enhancement